OpenStudy (anonymous):

A wave on a string, mass per unit length of 0.2 g/cm, travels in the z direction with the string displaced in the x and y directions as described by the equations x = 5 sin(kx - ωt) cm, and y = 5 cos(kx - ωt) cm. The wavelength and the wave velocity are λ=10 cm and v =9 cm/s respectively. What average power is transmitted along the string?

OpenStudy (anonymous):

 This is a picture of waves travelling in the z direction simplified into plane waves. I believe the formula I use here is P = I/a.

OpenStudy (vincent-lyon.fr):

If your wave travels in the z direction, then its equations must be: x = 5 sin(kz - ωt) cm y = 5 cos(kz - ωt) cm Your wave shows circular polarisation.

OpenStudy (vincent-lyon.fr):

Kinetic energy lineic density is: \(\large e_K=\frac 12 \mu v^2\), where v(z,t) is the velocity of a string's point. As displacement is 5 cm on a circle, velocity is 5ω, independent of z and t. With wave celerity and lineic mass of the string, you can find \(e_K=8.10^{-4}J.m^{-1}\)

OpenStudy (anonymous):

I'm still unsure how to compute the average power transmitted from here.

OpenStudy (anonymous):

also how can \[e_{K}\] come out with units J/m

OpenStudy (anonymous):

a) 0.2 × 10-4 W b) 0.1 × 10-3 W c) 0.4 × 10-5 W d) 1.4361×104 W

OpenStudy (vincent-lyon.fr):

I do not know what concepts you know and are able to use. It seems the easiest way is to computer lineic kinetic energy in the first place; Then lineic kinetic energy and elastic potential energy are in a ratio 1:1. So the overall lineic energy is \(\large e=2e_K=1.6\times 10^{-3}J.m^{-1}\) As energy is carried by the wave at the same celerity as the wave itself, power transmitted the product of lineic energy by celerity. Answer is P = e.v = 0.144 mW

OpenStudy (vincent-lyon.fr):

*is the product